A food contains 10 grams protein, 33 grams carbs, 6 grams of fat per serving. You consumed 1.5 servings of this food, what % of kilocalories from fat did you consume?
1
Expert's answer
2015-09-03T01:29:03-0400
A food contains 10 grams protein, 33 grams carbs, 6 grams of fat per serving. You consumed 1.5 servings of this food, what % of kilocalories from fat did you consume?

Given:
A food contains
m_p = 10 g
m_c = 33 g
m_f = 6 g
m_g = m_p + m_c + m_f = 49
m_c = 1.5*m_g = 73.5 g

Solution:
So in the food portion which mass 49 g
x_p = 10 g *100% /49 g = 20.41 % protein
x_c = 33 g *100% /49 g = 67.35 % carbs
x_f = 6 g *100% /49 g = 12.24 % fat

So in the food portion which mass 73.5 g

m_p_n = 20.41% * 73.5 / 100% = 15.00 g
m_c_n = 67.35% * 73.5 / 100% = 49.50 g
m_c_f = 12.24% * 73.5 / 100% = 9.00 g

Question: what % of kilocalories from fat did you consume?

Fat energy density 9 kcal/g
Proteints energy density 4 kcal/g
Carbc energy density 4 kcal/g


we have m_c_f = 9 g * 9 kcal/g = 81 kcal
m_c_n = 49.5 g * 4 kcal/g = 198 kcal
m_p_n = 15g *4 kcal/g = 60 kcal
Food General kcal = 81+198+60 = 339 kcal
So % fat = 81*100/339 = 23.89 %


Answer: 23.89 %